OpenStudy (anonymous):

Jason considered two similar televisions at a local electronics store. The generic version was based off the brand name and was 2/3 the size of the brand name. If the generic television set is 12 inches by 24 inches, what are the dimensions of the brand name television?

OpenStudy (wasiqss):

18 by 36

OpenStudy (anonymous):

How did you solve this?

OpenStudy (amistre64):

linear scales affect all linear measuers

OpenStudy (wasiqss):

like if generic version is 2/3 size of brand version , hence we divide the generic versin size by 2/3

OpenStudy (amistre64):

12*2/3 x 24*2/3

OpenStudy (anonymous):

But I did that ang got 8/16?

OpenStudy (amistre64):

i get 8 x 16

OpenStudy (anonymous):

But it is not 8 by 16

OpenStudy (paxpolaris):

No the brand version is bigger

hero (hero):

8 x 16 = 8 by 16

OpenStudy (amistre64):

then your question is bad

OpenStudy (amistre64):

or i read it wrong ....

OpenStudy (amistre64):

G(x,y) = 2/3 B(x,y) G(12,24) = 2/3 B(x,y)

OpenStudy (amistre64):

so, 12*3/2 x 24*3/2

hero (hero):

It doesn't really specify if the brand name was 2/3 bigger or 2/3 smaller

hero (hero):

By default, I would assume 2/3 smaller

OpenStudy (amistre64):

G = 2/3 B is specific enough tho

OpenStudy (anonymous):

you multiply both linear dimensions by 3/2 to get the dimensions of the brand name tv

OpenStudy (anonymous):

I know! Thats what I assumed as well.

OpenStudy (paxpolaris):

if L x W are the dimensions of the brand version:\[L \times \frac23=24\] \[W \times \frac23=12\]

OpenStudy (amistre64):

The dimensions given are the Generic; not the Brand

OpenStudy (anonymous):

correct, pax

hero (hero):

Yeah, I see now. I get it.

hero (hero):

I can see how someone could easily confuse the question though.

OpenStudy (anonymous):

So would I just multiply each value by 3/2?

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

Gotcha!

hero (hero):

Qwerty, the key is being able to read and understand the question

OpenStudy (anonymous):

Yes I understand that Hero.

hero (hero):

Jason considered two similar televisions at a local electronics store. The generic version was 2/3 the size of the brand name. If the generic television set is 12 inches by 24 inches, what are the dimensions of the brand name television? Here's the way the question should have been written if you get rid of the unnecessary info.

OpenStudy (anonymous):

Yes I got that. I know the generic television is 12 by 24 and I need to know the brand nme size if this is 2/3 of it.

OpenStudy (anonymous):

Thanks guys! I got a 100!

hero (hero):

good job